A salesperson of kathmandu mall has a 70% chance of making a sale to each customer. the behaviour of successive customer is assumed to be independent. if two customers A andB enter, what is the probability that the salesperson will make a sale to none and at least one?
Answer by ikleyn(52785)

.

(b)  The probability that the salesperson will make a sale to at least one


     P(make a sale to at least one) = P( (make a sale to one) OR ( make a sale to the second) ) = 0.7 + 0.7 - 0.7*0.7 

                                    = 1.4 - 0.49 = 0.91.    ANSWER



(a)  The probability that the salesperson will make a sale to none = complement to the probability of n. (b) 

                                    = 1 - 0.91 = 0.09.      ANSWER

Solved.